# ActAvatar: Temporally-Aware Precise Action Control for Talking Avatars

### v1

@@ 1 Introduction
Talking avatar generation has become increasingly important for applications ranging from virtual assistants [peng2023emotalk] and digital entertainment to online education [ye2023geneface] and telepresence systems [peng2025synctalk++]. Recent advances in diffusion models [wan2025wan, kong2024hunyuanvideo] have significantly improved the visual quality of generated talking avatars [cui2024hallo3, wei2025mocha, wang2025fantasytalking, peng2025omnisync, lin2025omnihuman, gan2025omniavatar, zhang2025uniavgen].
However, existing methods [zhang2023sadtalker, tu2025stableavatar, peng2024synctalk, lin2025omnihuman] face three critical limitations. First, while current models can generate plausible hand movements [zhou2025exges], they struggle to accurately execute specific actions described in prompts due to treating the entire prompt uniformly [kong2025let], where action-related descriptions compete with scene descriptions for attention. Second, actions may appear at arbitrary moments rather than synchronizing with semantically relevant speech segments [zhang2025uniavgen]. This temporal drift arises because standard conditioning mechanisms lack explicit temporal structure, causing attention to diffuse uniformly across time. Third, many methods resort to explicit control modalities such as pose skeleton sequences [meng2025echomimicv2, cui2025hallo4], which increase pipeline complexity and limit the ability to generate novel actions.
These limitations reveal a fundamental challenge: establishing explicit correspondences between language semantics (what actions to perform), temporal windows (when to perform them), and audio cues (how they relate to speech). Moreover, text-driven action generation and audio-driven lip synchronization represent competing objectives that can interfere during generation. When both modalities exert strong influence simultaneously, the model struggles to balance conflicting signals, often resulting in degraded action quality or compromised lip-sync accuracy [wang2020makes]. In addition, fine-tuning pre-trained models on domain-specific data to improve audio-visual alignment often leads to catastrophic forgetting, where the model’s original text-following capabilities are weakened or lost entirely.
@@ 2.1 Video Generation Models
Text-to-video generation [wu2023tune, singer2022make, rao2024modelgrow, lin2024open, zhang2025arbitrary, peng2024controlnext] has witnessed remarkable progress with the advent of diffusion models. Early works such as Video Diffusion Models [ho2022video] and Imagen Video [ho2022imagen] established foundational frameworks by extending image diffusion to the temporal domain through 3D U-Net architectures and factorized spatiotemporal attention. Subsequent methods like AnimateDiff [guo2023animatediff] and Stable Video Diffusion [blattmann2023stable] achieved substantial improvements in visual quality and motion coherence by leveraging pre-trained text-to-image models and introducing specialized temporal modeling modules.
Recently, video generation models have shifted towards Transformer-based [vaswani2017attention] architectures to achieve better scalability. HunyuanVideo [kong2024hunyuanvideo] creates a comprehensive training framework that advances efficient model training and inference, Wan [wan2025wan] builds a complete and open video generation model suite that drives the development of the open-source community, and SkyReels v2 [chen2025skyreels] further extends video generation duration to infinite-length generation.
@@ 2.2 Talking Avatar Generation
Talking avatar generation [cui2024hallo3, wang2024styletalk++, peng2025dualtalk, ma2025talkclip, meng2025echomimicv3, yang2025flowerdance, yang2025megadance, ma2023dreamtalk, peng2023selftalk, zhang2025morpheus, chen2025hunyuanvideo, zhou2024meta, hu2025ggtalker, tu2025stableavatar, wu2024vgg, ma2023styletalk, wang2024v, ma2025playmate2] aims to synthesize realistic human video with synchronized lip movements and natural expressions driven by audio input. EMO [tian2024emo] pioneered audio-conditioned video diffusion for talking heads, introducing audio cross-attention for lip synchronization. EchoMimic [chen2025echomimic]further improved controllability through multi-modal conditioning combining audio with visual references. Methods like Echomimic V2 [meng2025echomimicv2] and Hallo 4 [cui2025hallo4] resort to explicit pose guidance through skeleton sequences, introducing additional annotation requirements and limiting the naturalness of language-based interaction. MultiTalk [kong2025let] and HunyuanVideo-Avatar [chen2025hunyuanvideo] extend these approaches to natural motion generation, but suffer from poor text-following capability.
Recent works further optimize this task, such as Kling-Avatar [ding2025kling] and OmniHuman 1.5 [jiang2025omnihuman]. These methods rely on strong video generation models to achieve relatively good action generation capabilities, but still perform diffusion on global prompts, failing to achieve precise temporal action control. AgentAvatar [wang2023agentavatar] try to introduce a timeline method, but only generate facial expressions. Our method addresses this limitation by introducing structured temporal prompts that enable phase-level precision in action control through hierarchical prompt decomposition and phase-aware attention mechanisms.
@@ 3.1 Overview
ActAvatar aims to achieve precise temporal action control in talking avatar generation through structured prompt conditioning. Given an audio sequence $\mathbf{a}\in\mathbb{R}^{T_{a}\times D_{a}}$ and a reference image $\mathbf{I}_{\text{ref}}\in\mathbb{R}^{H\times W\times 3}$, our goal is to generate a video $\mathbf{V}=\{\mathbf{I}_{t}\}_{t=1}^{T}$ where the avatar exhibits accurate lip synchronization with the audio and performs specific actions at semantically appropriate temporal windows as described in the structured prompt $\mathbf{P}$. The structured prompt is automatically generated by a Multimodal Large Language Model (MLLM, e.g., Qwen3-Omni [xu2025qwen3]) based on the input image and audio content.
@@ 3.4.1 Stage 1: Audio Adapter Training and Extraction
Stage 1 establishes robust audio-visual correspondence by training on diverse and large-scale talking-head videos. We adopt the Flow Matching training paradigm [lipman2022flow]. Given the original latent representation $\mathbf{x}_{0}$ (data) and pure noise $\mathbf{x}_{1}\sim\mathcal{N}(0,\mathbf{I})$, we construct the optimal transport flow path:
@@
where $\mathbf{C}_{\text{brief}}$ represents brief text captions (e.g., “A woman speaking”), and $\mathbf{A}$ are audio embeddings from Wav2Vec 2.0 [baevski2020wav2vec]. The audio adapter consists of an audio projection module (mapping audio encodings to frame-aligned tokens) and audio cross-attention layers (fusing audio semantics via frame-wise attention).
@@ 3.4.2 Stage 2: Temporally-Aware Action Control
Stage 2 injects temporal action control through structured annotations. We construct a dataset via: (1) applying DWPose [yang2023effective] to compute motion magnitude and selecting videos with significant movement; (2) using a multimodal large language model to generate hierarchical prompts with base blocks and phase-specific descriptions with temporal anchors. This yields a focused dataset with high-quality structured annotations.
@@
Hallo3 [cui2024hallo3] 30.424 318.451 3.883 2.488 6.994 8.647 5B 480p 32 min
FantasyTalking [wang2025fantasytalking] 25.615 458.733 3.955 2.663 4.179 11.067 14B 720p 83 min
EchoMimic v3 [meng2025echomimicv3] 43.544 595.805 4.027 2.690 4.555 10.804 1.3B 480p 7 min
HunyuanVideo-Avatar [chen2025hunyuanvideo] 24.515 322.844 4.054 2.693 7.647 7.564 13B 720p 74 min
MultiTalk [kong2025let] 25.578 368.719 3.981 2.570 7.610 7.645 14B 480p 45 min
OmniAvatar [gan2025omniavatar] 24.398 374.594 4.088 2.664 7.986 7.696 14B 480p 36 min
StableAvatar [tu2025stableavatar] 23.742 323.601 3.984 2.572 5.457 10.151 1.3B 480p 12 min
Wan-S2V [gao2025wan] 23.850 299.491 4.108 2.684 7.462 7.745 14B 720p 68 min
@@
Hallo3 [cui2024hallo3] 4.288 10.266 4.462 3.531 0.756 5.299 6.541 6.793 7.844
FantasyTalking [wang2025fantasytalking] 4.209 10.948 4.510 3.560 0.728 4.995 6.435 6.501 8.147
EchoMimic v3 [meng2025echomimicv3] 3.199 11.896 4.301 3.543 0.764 4.717 6.684 5.817 6.115
HunyuanVideo-Avatar [chen2025hunyuanvideo] 6.251 8.862 4.759 3.725 0.674 3.977 5.491 6.609 8.044
MultiTalk [kong2025let] 6.654 8.356 4.446 3.477 0.574 3.490 4.654 5.455 7.584
OmniAvatar [gan2025omniavatar] 6.765 8.785 4.569 3.610 0.818 5.505 7.032 7.147 8.042
StableAvatar [tu2025stableavatar] 4.035 11.615 4.760 3.658 0.758 4.802 6.576 6.722 7.768
Wan-S2V [gao2025wan] 6.473 8.401 4.701 3.647 0.754 4.934 6.465 6.630 8.168
@@ 4.1.2 Datasets
Training Data. For Stage 1 audio-driven lip synchronization training, we utilize 500K diverse talking head videos from OpenHumanVid [li2025openhumanvid] and SpeakerVid [zhang2025speakervid], covering varied speakers, emotions, and speaking styles. For Stage 2, we construct a structured annotation dataset through DWPose-based motion selection and MLLM-based prompt generation, yielding 100K samples with phase-level temporal annotations. The detailed construction pipeline is provided in the supplementary material.
Evaluation Data. We evaluate on two test sets: (1) HDTF Test Set [zhang2021flow], containing 100 high-quality talking-head videos (5s each) focusing on lip synchronization quality, as it includes only upper body without hand movements; (2) Action Bench, our constructed benchmark consisting of 200 samples with diverse action instructions. Each sample includes a reference image, TTS-synthesized speech, and structured prompts with action annotations.The prompts are MLLM-generated with human verification.
@@ 4.1.3 Evaluation Metrics
We evaluate ActAvatar using comprehensive metrics. For lip sync, we use Sync-C and Sync-D from SyncNet [chung2016out] (higher Sync-C and lower Sync-D indicate better alignment). For visual quality, we report FID and FVD (lower is better), and use Q-Align [wu2023q] for video quality (IQA) and aesthetics (ASE) (higher is better). For action control, we develop a Gemini-based evaluation framework that provides: (1) Action Occurrence (AO): whether the described action appears; (2) Action Accuracy (AA): how well the action matches the description (0-10); (3) Temporal Correctness (TC): whether the action occurs in the specified time window (0-10); (4) Action Quality (AQ): overall naturalness of execution (0-10); (5) Hand Clarity (HC): hand quality (0-10). We report Hit@Segment (H@S) as the percentage of phases where AO=1, and mean scores for AA, TC, AQ, and HC. To ensure robustness, we run the Gemini evaluation 5 times for each video and report the average scores. The evaluation prompts and framework construction are provided in the supplementary material.
@@ 4.4 User Study
EchoMimic v3 [meng2025echomimicv3] 3.43 3.25 3.21 2.45 2.68
HunyuanVideo-Avatar [chen2025hunyuanvideo] 2.84 3.12 3.68 3.67 3.45
MultiTalk [kong2025let] 3.19 3.16 3.95 3.41 3.58
OmniAvatar [gan2025omniavatar] 3.75 3.92 4.08 3.56 3.61
Wan-S2V [gao2025wan] 3.67 3.78 4.12 3.68 4.04

### v2

@@ 1 Introduction
Talking avatar generation has become increasingly important for applications ranging from virtual assistants [35] and digital entertainment to online education [55] and telepresence systems [31]. Recent advances in diffusion models [41, 17] have significantly improved the visual quality of generated talking avatars [8, 47, 44, 33, 21, 10, 57].
However, existing methods [58, 39, 32, 21] face three critical limitations. First, while current models can generate plausible hand movements [62], they struggle to accurately execute specific actions described in prompts due to treating the entire prompt uniformly [18], where action-related descriptions compete with scene descriptions for attention. Second, actions may appear at arbitrary moments rather than synchronizing with semantically relevant speech segments [57]. This temporal drift arises because standard conditioning mechanisms lack explicit temporal structure, causing attention to diffuse uniformly across time. Third, many methods resort to explicit control modalities such as pose skeleton sequences [28, 7], which increase pipeline complexity and limit the ability to generate novel actions.
These limitations reveal a fundamental challenge: establishing explicit correspondences between language semantics (what actions to perform), temporal windows (when to perform them), and audio cues (how they relate to speech). Moreover, text-driven action generation and audio-driven lip synchronization represent competing objectives that can interfere during generation. When both modalities exert strong influence simultaneously, the model struggles to balance conflicting signals, often resulting in degraded action quality or compromised lip-sync accuracy [46]. In addition, fine-tuning pre-trained models on domain-specific data to improve audio-visual alignment often leads to catastrophic forgetting, where the model’s original text-following capabilities are weakened or lost entirely.
@@ 2.1 Video Generation Models
Text-to-video generation [50, 37, 36, 20, 56, 29] has witnessed remarkable progress with the advent of diffusion models. Early works such as Video Diffusion Models [14] and Imagen Video [13] established foundational frameworks by extending image diffusion to the temporal domain through 3D U-Net architectures and factorized spatiotemporal attention. Subsequent methods like AnimateDiff [12] and Stable Video Diffusion [2] achieved substantial improvements in visual quality and motion coherence by leveraging pre-trained text-to-image models and introducing specialized temporal modeling modules.
Recently, video generation models have shifted towards Transformer-based [40] architectures to achieve better scalability. HunyuanVideo [17] creates a comprehensive training framework that advances efficient model training and inference, Wan [41] builds a complete and open video generation model suite that drives the development of the open-source community, and SkyReels v2 [3] further extends video generation duration to infinite-length generation.
@@ 2.2 Talking Avatar Generation
Talking avatar generation [8, 45, 30, 24, 27, 53, 52, 26, 34, 61, 4, 63, 15, 39, 49, 25, 42, 23] aims to synthesize realistic human video with synchronized lip movements and natural expressions driven by audio input. EMO [38] pioneered audio-conditioned video diffusion for talking heads, introducing audio cross-attention for lip synchronization. EchoMimic [5]further improved controllability through multi-modal conditioning combining audio with visual references. Methods like Echomimic V2 [28] and Hallo 4 [7] resort to explicit pose guidance through skeleton sequences, introducing additional annotation requirements and limiting the naturalness of language-based interaction. MultiTalk [18] and HunyuanVideo-Avatar [4] extend these approaches to natural motion generation, but suffer from poor text-following capability.
Recent works further optimize this task, such as Kling-Avatar [9] and OmniHuman 1.5 [16]. These methods rely on strong video generation models to achieve relatively good action generation capabilities, but still perform diffusion on global prompts, failing to achieve precise temporal action control. AgentAvatar [43] try to introduce a timeline method, but only generate facial expressions. Our method addresses this limitation by introducing structured temporal prompts that enable phase-level precision in action control through hierarchical prompt decomposition and phase-aware attention mechanisms.
@@ 3.1 Overview
ActAvatar aims to achieve precise temporal action control in talking avatar generation through structured prompt conditioning. Given an audio sequence $\mathbf{a}\in\mathbb{R}^{T_{a}\times D_{a}}$ and a reference image $\mathbf{I}_{\text{ref}}\in\mathbb{R}^{H\times W\times 3}$, our goal is to generate a video $\mathbf{V}=\{\mathbf{I}_{t}\}_{t=1}^{T}$ where the avatar exhibits accurate lip synchronization with the audio and performs specific actions at semantically appropriate temporal windows as described in the structured prompt $\mathbf{P}$. The structured prompt is automatically generated by a Multimodal Large Language Model (MLLM, e.g., Qwen3-Omni [51]) based on the input image and audio content.
@@ 3.4.1 Stage 1: Audio Adapter Training and Extraction
Stage 1 establishes robust audio-visual correspondence by training on diverse and large-scale talking-head videos. We adopt the Flow Matching training paradigm [22]. Given the original latent representation $\mathbf{x}_{0}$ (data) and pure noise $\mathbf{x}_{1}\sim\mathcal{N}(0,\mathbf{I})$, we construct the optimal transport flow path:
@@
where $\mathbf{C}_{\text{brief}}$ represents brief text captions (e.g., “A woman speaking”), and $\mathbf{A}$ are audio embeddings from Wav2Vec 2.0 [1]. The audio adapter consists of an audio projection module (mapping audio encodings to frame-aligned tokens) and audio cross-attention layers (fusing audio semantics via frame-wise attention).
@@ 3.4.2 Stage 2: Temporally-Aware Action Control
Stage 2 injects temporal action control through structured annotations. We construct a dataset via: (1) applying DWPose [54] to compute motion magnitude and selecting videos with significant movement; (2) using a multimodal large language model to generate hierarchical prompts with base blocks and phase-specific descriptions with temporal anchors. This yields a focused dataset with high-quality structured annotations.
@@
Hallo3 [8] 30.424 318.451 3.883 2.488 6.994 8.647 5B 480p 32 min
FantasyTalking [44] 25.615 458.733 3.955 2.663 4.179 11.067 14B 720p 83 min
EchoMimic v3 [27] 43.544 595.805 4.027 2.690 4.555 10.804 1.3B 480p 7 min
HunyuanVideo-Avatar [4] 24.515 322.844 4.054 2.693 7.647 7.564 13B 720p 74 min
MultiTalk [18] 25.578 368.719 3.981 2.570 7.610 7.645 14B 480p 45 min
OmniAvatar [10] 24.398 374.594 4.088 2.664 7.986 7.696 14B 480p 36 min
StableAvatar [39] 23.742 323.601 3.984 2.572 5.457 10.151 1.3B 480p 12 min
Wan-S2V [11] 23.850 299.491 4.108 2.684 7.462 7.745 14B 720p 68 min
@@
Hallo3 [8] 4.288 10.266 4.462 3.531 0.756 5.299 6.541 6.793 7.844
FantasyTalking [44] 4.209 10.948 4.510 3.560 0.728 4.995 6.435 6.501 8.147
EchoMimic v3 [27] 3.199 11.896 4.301 3.543 0.764 4.717 6.684 5.817 6.115
HunyuanVideo-Avatar [4] 6.251 8.862 4.759 3.725 0.674 3.977 5.491 6.609 8.044
MultiTalk [18] 6.654 8.356 4.446 3.477 0.574 3.490 4.654 5.455 7.584
OmniAvatar [10] 6.765 8.785 4.569 3.610 0.818 5.505 7.032 7.147 8.042
StableAvatar [39] 4.035 11.615 4.760 3.658 0.758 4.802 6.576 6.722 7.768
Wan-S2V [11] 6.473 8.401 4.701 3.647 0.754 4.934 6.465 6.630 8.168
@@ 4.1.2 Datasets
Training Data. For Stage 1 audio-driven lip synchronization training, we utilize 500K diverse talking head videos from OpenHumanVid [19] and SpeakerVid [59], covering varied speakers, emotions, and speaking styles. For Stage 2, we construct a structured annotation dataset through DWPose-based motion selection and MLLM-based prompt generation, yielding 100K samples with phase-level temporal annotations. The detailed construction pipeline is provided in the supplementary material.
Evaluation Data. We evaluate on two test sets: (1) HDTF Test Set [60], containing 100 high-quality talking-head videos (5s each) focusing on lip synchronization quality, as it includes only upper body without hand movements; (2) Action Bench, our constructed benchmark consisting of 200 samples with diverse action instructions. Each sample includes a reference image, TTS-synthesized speech, and structured prompts with action annotations.The prompts are MLLM-generated with human verification.
@@ 4.1.3 Evaluation Metrics
We evaluate ActAvatar using comprehensive metrics. For lip sync, we use Sync-C and Sync-D from SyncNet [6] (higher Sync-C and lower Sync-D indicate better alignment). For visual quality, we report FID and FVD (lower is better), and use Q-Align [48] for video quality (IQA) and aesthetics (ASE) (higher is better). For action control, we develop a Gemini-based evaluation framework that provides: (1) Action Occurrence (AO): whether the described action appears; (2) Action Accuracy (AA): how well the action matches the description (0-10); (3) Temporal Correctness (TC): whether the action occurs in the specified time window (0-10); (4) Action Quality (AQ): overall naturalness of execution (0-10); (5) Hand Clarity (HC): hand quality (0-10). We report Hit@Segment (H@S) as the percentage of phases where AO=1, and mean scores for AA, TC, AQ, and HC. To ensure robustness, we run the Gemini evaluation 5 times for each video and report the average scores. The evaluation prompts and framework construction are provided in the supplementary material.
@@ 4.4 User Study
EchoMimic v3 [27] 3.43 3.25 3.21 2.45 2.68
HunyuanVideo-Avatar [4] 2.84 3.12 3.68 3.67 3.45
MultiTalk [18] 3.19 3.16 3.95 3.41 3.58
OmniAvatar [10] 3.75 3.92 4.08 3.56 3.61
Wan-S2V [11] 3.67 3.78 4.12 3.68 4.04